# Complexity of deep computations via topology of function spaces

### v2

@@ Introduction
In this paper we study asymptotic behavior of computations, e.g., the depth of a neural network tending to infinity, or the time interval between layers of a time-series network tending toward zero. Recently, particular cases of this concept have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [8], Physics-Informed Neural Networks [RPK19], and deep equilibrium models [4]). The formal framework introduced here provides a unified setting to study these limit phenomena from a foundational viewpoint.
@@
In a previous paper [ADI+24], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
@@
In [ADI+24], we proved a new “tame vs wild” (i.e., polynomially approximable vs nonpolynimially approximable) dichotomy for the complexity of deep computations by invoking a classical result of Grothendieck from the 1950s [GRO52]. Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in model theory.
@@
Simplest among deep computations are those arising as pointwise limits of (continuous) computations proper. In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions. Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity away from the Baire class 0 which (by definition) consists of continuous functions.
@@
We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
@@
Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways; since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22]. It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory. At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
@@
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [VAL84].
@@
Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; they identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes. They also showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
@@
We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields. In the interest of accessibility, we do not assume the reader to have previous familiarity with advanced topology, model theory, or computing. The only technical prerequisites to read this paper are undergraduate-level topology and measure theory. The necessary topological background beyond undergraduate topology is covered in section 1.
@@ 1. General topological preliminaries: From continuity to Baire class 1
In this section we present some preliminaries from general topology and function space theory. In the interest of completeness, we some proofs that may be safely skipped by readers familiar with these topics.
@@
A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions. If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$). Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)$. The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
@@
A topological space $X$ is Baire if every countable intersection of dense open sets is dense. The Baire Category Theorem states that every Hausdorff compact or completely metrizable space (hence, every Polish space) is Baire.
The following fact was established by Baire in his 1899 thesis. A proof can be found in [TOD97, Section 10].
@@
    $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ whenever $U\subseteq Y$ is open.
If, mooreover, $X$ is Baire, then (1) and (2) are equivalent to:
@@
A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact. Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest analysis and topological dynamics. We begin with the following well-known result. Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$. We include a proof for the reader’s convenience:
@@
    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has an accumulation point in $B_{1}(X)$.
@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, points of the closure of some subset can always be approximated by points in the set proper, via a convergent sequence. For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways. The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ is closed, but has no convergent subsequences, in $Y=C_{p}(X)$: in a sense, such example exhibits the worst failure of sequential convergence possible. The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) realizes the Stone-Čech compactification $\beta\mathbb{N}$ of the discrete space of natural numbers; it is an important object of study in general topology.
The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach spaces.
@@
Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP. Reciprocally, assume that $X$ is compact and has finitary IP. Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$. For any $I,J\subseteq\mathbb{N}$ (eventually disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$. For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively. Clearly, $E,F\subseteq\mathbb{N}$ witness finitary NIP for $\{f_{n}\}$ if and only if $X^{E,F}=\emptyset$.
@@
Let $X$ be a metrizable compact (hence Polish) space. For every pointwise bounded $A\subseteq C_{p}(X)$, the following properties are all equivalent:
@@ 1.2. NIP as a universal dividing line between polynomial and exponential complexity
$\sum_{i=0}^{N-1}\binom{m}{i}\leq\sum_{i=0}^{N-1}\frac{m^{i}}{i!}\leq\left(\sum_{i=0}^{N-1}\frac{1}{i!}\right)\cdot m^{N-1}\cdot<e\cdot m^{N-1}=O\left(m^{N}\right).$
@@ 1.3. Rosenthal compacta
The identification of $C_{p}$-spaces follows trivially from the definition of topological product and the fact that $\mathcal{P}$ is discrete: a continuous map $X\to\mathbb{R}^{\mathcal{P}}$ is precisely a $\mathcal{P}$-tuple of continuous functions $X\to\mathbb{R}$, and these correspond to continuous functions $X\times\mathcal{P}\to\mathbb{R}$. The identification of Baire-1 spaces follows immediately, since it is defined in terms of the purely topological notion of limit (in the ambient space) of sequences of continuous functions. ∎
@@
2. Compositional Computation Structures: A structural approach to floating-point computations
In this section, we connect function spaces with floating point computations. We start by summarizing some basic concepts from [ADI+24].
@@ 2. Compositional Computation Structures: A structural approach to floating-point computations
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$. Important space states are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$. We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$. Via the identification $v\mapsto\operatorname{tp}(v)$, the states space $L$ is correspondingly topologized; in particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
@@
Intuitively, state types capture a notion of “ultrastate”.
The traditional setting of discrete first-order logic may be regarded as the special case in which predicates are binary (such predicates may take only the values 0 and $1$, say). In such a traditional setting, the space of types of a structure is necessarily compact; for some purposes, it serves as a sort of compactification of the structure itself. Compactness in such setting is critical. However, predicates of a computational states structure take values in the non-compact space $\mathbb{R}$, so the corresponding type space $\mathcal{L}$ need not be compact. To bypass this obstacle, we follow the idea, introduced in [ADI+24], of covering $\mathcal{L}$ by “thin” compact subspaces called shards.
@@
A sizer is a tuple $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
@@
Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards (as the sizer $r_{\bullet}$ varies).
In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$ may be a proper subset. Equality holds in the important special case when $\mathcal{P}$ is countable (see *** below).

### v3

@@ Introduction
In this paper we study asymptotic behavior of computations as parameters of the computation tend towards a limit, e.g., the depth of a neural network tending to infinity, or the time interval between layers of a time-series network tending toward zero. Recently, particular cases of this concept have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [8], Physics-Informed Neural Networks [RPK19], and deep equilibrium models [4]). The formal framework introduced here provides a unified setting to study these limit phenomena from a foundational viewpoint.
@@
In a previous paper [ADI+24], the authors introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
@@
In [ADI+24], a new “tame vs wild” (i.e., polynomially approximable vs nonpolynomially approximable) dichotomy for the complexity of deep computations was proved by invoking a classical result of Grothendieck from the 1950s [GRO52]. Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in the sense of model theory.
@@
Deep computations arise as topological limits of standard (continuous) computations. Simplest among these is when the limit is sequential. In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions. Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity above the Baire class 0 which (by definition) consists of continuous functions.
@@
In this paper, we apply the theory of spaces of Baire class 1 functions to identify and separate new complexity classes of deep computations. We obtain a new “tame vs wild” dichotomy for deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and further, a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
@@
Todorčević’s trichotomy concerns Rosenthal compacta, which are topological spaces that fit our context superbly: they are defined as compact spaces that are homeomorphic to subspaces of the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways. Since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy for Rosenthal compacta has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22]. It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory. At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
@@
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of Non Independence Property (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concepts VC-dimension and Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory [VC71, VC74, VAL84].
@@
Refining Todorčević’s trichotomy, we invoke a more recent heptachotomy for separable Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; these authors identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
@@
We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields. In the interest of accessibility, we do not assume that the reader has previous familiarity with advanced topology, model theory, or computing. The only technical prerequisites to read this paper are undergraduate-level topology and measure theory. The necessary topological background beyond undergraduate topology is covered in section 1.
@@ 1. General topological preliminaries: From continuity to Baire class 1
In this section we present some preliminaries from general topology and function space theory. In the interest of completeness, we include some proofs that may be safely skipped by readers familiar with these topics.
@@
A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions. In symbols, $f:X\to Y$ is Baire class 1 if there is a sequence of continuous functions $f_{n}:X\to Y$ such that for all $x\in X$, $f(x)=\lim_{n\to\infty}f_{n}(x)$. If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$). Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)\subseteq Y^{X}$ and we give these the topology (called the topology of pointwise convergence) inherited from the product topology of $Y^{X}$. The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits. An elementary fact about Baire class 1 functions is that they are continuous except on a set of first category (also called a meager set, a set of first cateogory is the countable union of sets whose closure has empty interior; intuitively, these sets are “topologically small”). Thus, Baire class 1 functions are continuous on a “topologically large” subset of their domain.
@@
A topological space $X$ is Baire if every countable intersection of dense open sets is dense. The Baire Category Theorem states that every compact Hausdorff or completely metrizable space (hence, every Polish space) is Baire.
The following fact was established by Baire in his groundbreaking thesis. A proof can be found in [TOD97, Section 10].
@@
    $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ whenever $U\subseteq\mathbb{R}$ is open.
If, moreover, $X$ is Baire, then (1) and (2) are equivalent to:
@@
A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact. Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest in analysis and topological dynamics. We begin with the following well-known result. Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$. We include a proof for the reader’s convenience:
@@
    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has a limit point in $B_{1}(X)$.
@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, points of the closure of some subset can always be approximated by points in the set proper, via a convergent sequence. For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways. The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ has no convergent subsequences, in $\mathbb{R}^{X}$. In a sense, this example exhibits the worst failure of sequential convergence possible. The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) is homeomorphic to the Stone-Čech compactification of the discrete space of natural numbers, usually denoted $\beta\mathbb{N}$, which is an important object of study in general topology.
The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach space.
@@
Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP. Reciprocally, assume that $X$ is compact and has finitary IP. Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$. For any $I,J\subseteq\mathbb{N}$ (almost disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$. For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively. Clearly, $E,F\subseteq\mathbb{N}$ witness finitary NIP for $\{f_{n}\}$ if and only if $X^{E,F}=\emptyset$.
@@
Let $X$ be a compact metrizable (hence Polish) space. For every pointwise bounded $A\subseteq C_{p}(X)$, the following properties are all equivalent:
@@ 1.2. NIP as a universal dividing line between polynomial and exponential complexity
$\sum_{i=0}^{N-1}\binom{m}{i}\leq\sum_{i=0}^{N-1}\frac{m^{i}}{i!}\leq\left(\sum_{i=0}^{N-1}\frac{1}{i!}\right)\cdot m^{N-1}<e\cdot m^{N-1}=O\left(m^{N}\right).$
@@ 1.3. Rosenthal compacta
The identification of $C_{p}$-spaces follows trivially from the definition of topological product and the fact that $\mathcal{P}$ is discrete: a continuous map $X\to\mathbb{R}^{\mathcal{P}}$ is precisely a $\mathcal{P}$-indexed family of continuous functions $X\to\mathbb{R}$, and these correspond to continuous functions $X\times\mathcal{P}\to\mathbb{R}$. The identification of Baire-1 spaces follows immediately, since it is defined in terms of the purely topological notion of limit (in the ambient space) of sequences of continuous functions. ∎
@@
2. Compositional Computation Structures: A structural approach to arbitrary-precision arithmetic
In this section, we connect function spaces with arbitrary-precision arithmetic computations. We start by summarizing some basic concepts from [ADI+24].
@@ 2. Compositional Computation Structures: A structural approach to arbitrary-precision arithmetic
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$. Important state spaces are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$. We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$. Via the identification $v\mapsto\operatorname{tp}(v)$, the states space $L$ is correspondingly topologized; in particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
@@
Intuitively, state types capture a notion of “limit state”.
As we combine ideas of model theory [KEI03] and topology [BFT78], we are interested in families of real-valued functions that are pointwise bounded. This leads us to the concepts of sizer and shard introduced first in introduced in [ADI+24]:
@@
A sizer is a family $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
@@
Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards as the sizer $r_{\bullet}$ varies.
In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$, and the inclusion may be proper. However, equality holds in the important special case when $\mathcal{P}$ is countable (see [ADI+24]).